# Neural Implicit Flow Fields for Spatio-Temporal Motion Mapping

### v1

@@ Abstract
Safe and efficient robot operation in complex human environments can benefit from good models of site-specific motion patterns. Maps of Dynamics (MoDs) provide such models by encoding statistical motion patterns in a map, but existing representations use discrete spatial sampling and typically require costly offline construction. We propose a continuous spatio-temporal MoD representation based on implicit neural functions that directly map coordinates to the parameters of a Semi-Wrapped Gaussian Mixture Model. This removes the need for discretization and imputation for unevenly sampled regions, enabling smooth generalization across both space and time. Evaluated on a large public dataset with long-term real-world people tracking data, our method achieves better accuracy of motion representation and smoother velocity distributions in sparse regions while still being computationally efficient, compared to available baselines. The proposed approach demonstrates a powerful and efficient way of modeling complex human motion patterns.
@@ 1 Introduction
As illustrated in Fig. 1, incorporating MoDs into motion planning provides benefits in crowded environments, since they encode information about the expected motion outside of the robot’s sensor range, allowing for less reactive behavior. In the example shown in Fig. 1, the oncoming pedestrian flow is initially outside the robot’s observation radius. Without MoD awareness, the robot chooses a direct path to the goal but later becomes trapped in the oncoming crowd. In contrast, a planner informed by MoDs can exploit prior knowledge of human motion patterns to generate a trajectory that aligns with the expected flow, allowing the robot to reach the goal safely and efficiently. MoDs can also be applied to long-term human motion prediction (Zhu et al., 2023). As shown in the right of Fig. 1, MoDs help predict realistic trajectories that implicitly respect the complex topology of the environment, such as navigating around corners or avoiding obstacles.
@@
However, existing MoD representations require spatial discretization, with a manually selected map resolution for point locations and interpolation to estimate motion at arbitrary positions. This discretization introduces information loss, reduces flexibility, and complicates tuning across different environments.
@@
We evaluate our approach on real-world datasets of human motion and show that continuous MoDs not only improve representation accuracy but can also drastically reduce map construction time. Our method yields smoother and more consistent velocity distributions, resulting in more accurate representations of human motion patterns. In contrast to baseline approaches that rely on time-consuming per-cell motion modeling, it computes nearly two orders of magnitude faster than CLiFF-map Kucner et al. (2017). Unlike the faster but discretised representation of STeF-map Molina et al. (2022), our method preserves non-discretised directions, yielding results closer in spirit to CLiFF.
@@
Figure 1: Indicative applications of Maps of dynamics (MoD) for motion planning and human motion prediction. Left: Illustration of how MoDs can support socially aware motion planning. The robot (red diamond) navigates toward the goal (green cross) in the presence of two opposing human flows, with the underlying MoD shown as colored arrows. In this scenario, the oncoming pedestrian flow moving in the opposing direction is initially outside the robot’s observation radius (grey dashed circle). Without guidance from the MoD, the planner initially takes a direct path to the goal but eventually becomes stuck and collides with the oncoming flow. In contrast, when informed by the MoD, the robot aligns its trajectory with the motion patterns and reaches the goal efficiently and safely. Right: Human motion prediction with a $60\text{\,}\mathrm{s}$ horizon. The red line represents the ground truth trajectory and the green line represents the observed trajectory. With MoD guidance, CLiFF-LHMP Zhu et al. (2023) makes more accurate and realistic predictions than deep learning methods. While the trajectories predicted by Social LSTM (Alahi et al., 2016), TUTR (Shi et al., 2023) and MID (Gu et al., 2022) often are unfeasible (e.g., crossing the walls), CLiFF-LHMP predictions implicitly follow the topology of the environment.
@@ 2 Related Work
There are several types of MoDs described in the literature, generally striving to provide an efficient tool for storing and querying information about historical or expected changes in states within the environment. Occupancy-based methods focus on mapping human dynamics on occupancy grid maps, modeling motion as shifts in occupancy (Wang et al., 2015; 2016). Trajectory-based methods extract human trajectories and group them into clusters, with each cluster representing a typical path through the environment (Bennewitz et al., 2005). These approaches suffer from noisy or incomplete trajectories. To address this, Chen et al. (2016) formulate trajectory modeling as a dictionary learning problem and use augmented semi-nonnegative sparse coding to find local motion patterns characterized by partial trajectory segments.
MoDs can also be based on velocity observations. With velocity mapping, human dynamics can be modeled through flow models. Kucner et al. (2017) presented a probabilistic framework for mapping velocity observations, which is named Circular-Linear Flow Field map (CLiFF-map). CLiFF-map represents local flow patterns as a multi-modal, continuous joint distribution of speed and orientation, as further described in  Sec. 3. A benefit of CLiFF-map is that it can be built from incomplete or spatially sparse velocity observations (de Almeida et al., 2024), without the need to store a long history of data or deploy advanced tracking algorithms. CLiFF-maps are typically built offline, for the reason of high computational costs associated with the building process. This constraint limits their applicability in real environments.
@@ 3.2 Learning Continuous Motion Fields
In many real-world environments, human motion patterns exhibit strong daily periodicity, such as morning and evening rush hours or lunchtime activity. Motivated by this structure, we model time as a periodic variable and condition the MoD on the time of day. This assumption allows the representation to capture long-term temporal variations without requiring sequential rollouts, and enables efficient queries of motion dynamics at arbitrary spatio-temporal coordinates $(x,y,t)$.
@@
where $\mathbf{x}_{i}\in\mathbb{R}^{2}$ is the spatial coordinate, $t_{i}\in[0,1]$ is the normalized time of a day, and $\mathbf{v}_{i}=[\rho_{i},\theta_{i}]^{\top}$ is observed velocity, we learn a continuous function $\Phi_{\theta}$ that maps a spatio-temporal coordinate $(\mathbf{x},t)$ to SWGMM parameters:
@@
$p(\mathbf{v}_{i}\mid\Phi_{\theta}(\mathbf{x}_{i},t_{i}))=\sum_{j=1}^{J}w_{j}(\mathbf{x}_{i},t_{i})\,\mathcal{N}^{\mathrm{SW}}_{\bm{\mu}_{j}(\mathbf{x}_{i},t_{i}),\,\bm{\Sigma}_{j}(\mathbf{x}_{i},t_{i})}(\mathbf{v}_{i}),$ (4)
@@
$\mathcal{L}(\theta)=-\frac{1}{N}\sum_{i=1}^{N}\log\,p\big(\mathbf{v}_{i}\mid\Phi_{\theta}(\mathbf{x}_{i},t_{i})\big).$ (5)
@@ 4 Results
4.1 Dataset
[…]
@@ 4.2 Baselines
Circular-Linear Flow Field Map (CLiFF-map)
@@
Spatio-Temporal Flow Map (STeF-map)
@@
Online CLiFF-map
Online CLiFF-map (Zhu et al. (2025)) extends the static CLiFF model by updating the SWGMM parameters incrementally as new motion observations become available. Each grid location maintains an SWGMM, which is initialized upon first receiving observations and subsequently updated using the stochastic expectation-maximization (sEM) algorithm (Cappé & Moulines (2009)). In sEM, the expectation step of the original EM algorithm is replaced by a stochastic approximation step, while the maximization step remains unchanged. Like the static CLiFF-map, Online CLiFF-map outputs SWGMM parameters at each grid location, but supports continuous adaptation over time. In the experiments, we follow a spatio-temporal setting by generating an online CLiFF-map for each hour. Observations collected in an hour interval are treated as the new data batch for updating the SWGMMs, producing a temporally adaptive representation of motion dynamics.
4.3 Implementation Details
The output of the network $\Phi_{\theta}$ parameterizes an SWGMM over speed and orientation. For $J$ mixture components, the network predicts $6J$ raw values per query coordinate. Each component $j$ is defined by: a mixture weight $w_{j}$, obtained by applying a softmax over the raw weights; a mean speed $\mu_{j,s}=\max(0,\,\tilde{\mu}_{j,s})$ and mean orientation $\mu_{j,a}=\tilde{\mu}_{j,a}\bmod 2\pi$; variances $\sigma^{2}_{j,s}=\exp(\mathrm{clamp}(\tilde{v}_{j,s},-10,10))$ and $\sigma^{2}_{j,a}=\exp(\mathrm{clamp}(\tilde{v}_{j,a},-10,10))$; and a correlation coefficient $\rho_{j}=0.99\tanh(\tilde{\rho}_{j})$. Altogether, the network defines a valid SWGMM with parameters as in Eq. 3, where $\bm{\mu}_{j}=\big(\mu_{j,s},\,\mu_{j,a}\big)$ and $\bm{\Sigma}_{j}$ is the covariance matrix with diagonal entries $\sigma^{2}_{j,s},\,\sigma^{2}_{j,a}$ and correlation $\rho_{j}$. In the experiments, $J$ is set to 3 and coordinates are normalized to [1,1][-1,1]. Spatial input is processed by an MLP with hidden sizes [128, 64][128,\,64] and ReLU activations. Temporal input is processed by a two-layer SIREN (sine activations with $\omega_{0}^{(1)}=30$ in the first layer and $\omega_{0}^{(h)}=1$ in the hidden layer). The two streams are fused via FiLM modulation (Perez et al., 2018). The fused representation is passed to a linear head producing $6J$ outputs. Models are trained using the Adam optimizer with learning rate $10^{-3}$ for 100 epochs. An ablation of alternative temporal encodings is provided in Sec. 4.6.
@@ 4.4 Quantitative Results
To quantitatively evaluate the accuracy of modeling human motion patterns (MoD quality), we use the negative log-likelihood (NLL). An MoD represents human motion as a probability distribution over velocity conditioned on a spatio-temporal coordinate $(x,y,t)$, implemented as either an SWGMM (our method and CLiFF-maps) or a histogram (STeF-map). To evaluate representation accuracy, we use test data consisting of observed human motions in the same environment. For each test sample $(x,y,t)$, we query the MoD to obtain the predicted distribution and compute the likelihood of the observed motion under this distribution. A higher likelihood indicates that the predicted distribution better aligns with the observed data. We report NLL for numerical stability and easy comparison, so lower NLL values correspond to more accurate motion representations, i.e., higher quality MoDs.
Table 1 reports the accuracy results. Our method achieves the lowest NLL ($0.775\pm 2.052$), outperforming all baselines. Online CLiFF-map, CLiFF-map, and STeF-map exhibit significantly higher NLLs, with paired t-tests showing $p<0.001$ under the null hypothesis that baseline NLL is less than or equal to ours. The reductions relative to our method are respectively +0.752+0.752 (online CLiFF), +1.189+1.189 (CLiFF), and +4.801+4.801 (STeF), all with 95% confidence intervals.
@@
We also compare the map building time of the baselines against our approach as shown in Table 2. For the baselines, the training time corresponds to convergence on all grid cells, while for our method it corresponds to the neural network training time. Our method trains in 19 minutes, substantially faster than CLiFF-map (over 30 hours) while achieving higher accuracy. These results highlight the practicality of continuous MoDs for real-time applications, combining both accuracy and efficiency.
@@ 4.5 Qualitative Results
Examples of NeMo-map are shown in Fig. 5. The model is queried at regular spatial intervals at three different times of day, at locations where human motion appears in the training dataset. Across the day, the map adapts smoothly to changes in human motion patterns. For example, in the east corridor (right side of the ATC map), the flow is directed left/upwards in the morning, shifts direction at noon, where pedestrians keep left when facing oncoming flows, and turns into right/downwards in the evening. (These patterns are most clearly seen when displaying only the SWGMM mixture component with the largest weight, in the bottom row, but please note that the map maintains a representation of the full multimodal distribution at all times.) The generated flow fields capture such temporal variations and implicitly align with the environment’s topology, even though no explicit map was provided during training. For instance, speeds decrease near resting benches, motion flows pass through exits, and flows follow the corridors.
09:00 11:00 18:00
Refer to captionAll Refer to caption Refer to caption Refer to caption
Refer to captionMax Refer to caption Refer to caption Refer to caption
Figure 5: NeMo-map in the ATC dataset, for 09:00 (left), 11:00 (middle) and 18:00 (right), showing changes of motion patterns throughout the day. Predicted Semi-Wrapped Gaussian Mixture Models (SWGMMs) are visualized. At each location, arrow color encodes orientation and arrow length encodes speed. The top row shows multimodality by rendering all SWGMM components with transparency proportional to their weights, while the bottom row more clearly shows the dominant flow, only displaying the mixture component with the largest weight.

### v2

# NeMo-map: Neural Implicit Flow Fields for Spatio-Temporal Motion Mapping

@@ Abstract
Safe and efficient robot operation in complex human environments can benefit from good models of site-specific motion patterns. Maps of Dynamics (MoDs) provide such models by encoding statistical motion patterns in a map, but existing representations use discrete spatial sampling and typically require costly offline construction. We propose a continuous spatio-temporal MoD representation based on implicit neural functions that directly map coordinates to the parameters of a Semi-Wrapped Gaussian Mixture Model. This removes the need for discretization and imputation for unevenly sampled regions, enabling smooth generalization across both space and time. Evaluated on two public datasets with real-world people tracking data, our method achieves better accuracy of motion representation and smoother velocity distributions in sparse regions while still being computationally efficient, compared to available baselines. The proposed approach demonstrates a powerful and efficient way of modeling complex human motion patterns and high performance in the trajectory prediction downstream task.
@@ 1 Introduction
Incorporating MoDs into motion planning provides benefits in crowded environments, since they encode information about the expected motion outside of the robot’s sensor range, allowing for less reactive behavior. MoDs can also be applied to long-term human motion prediction (Zhu et al., 2023). MoDs help predict realistic trajectories that implicitly respect the complex topology of the environment, such as navigating around corners or avoiding obstacles.
@@
However, existing MoDs require spatial discretization, with a manually selected map resolution for point locations and interpolation to estimate motion at arbitrary positions. This discretization introduces information loss, reduces flexibility, and complicates tuning across different environments.
@@
We evaluate our approach on real-world datasets and show that continuous MoDs not only improve representation accuracy but can also be computationally efficient. Our method yields smoother and more consistent velocity distributions, resulting in more accurate representations of human motion patterns and higher performance in the trajectory prediction downstream task. In contrast to baseline approaches that rely on time-consuming per-cell motion modeling, it trains a model that represents motion continuously over space and time and enables fast inference at arbitrary locations. Unlike the faster but discretised representation of STeF-map Molina et al. (2022), our method preserves non-discretised directions, yielding results closer in spirit to CLiFF.
@@ 2 Related Work
Several MoDs types are described in the literature, providing an efficient tool for storing and querying about historical or expected changes in environment states. Occupancy-based methods focus on mapping dynamics on occupancy grids, modeling motion as shifts in occupancy (Wang et al., 2015; 2016). Trajectory-based methods extract trajectories and group into clusters, with each cluster representing a typical path through the environment (Bennewitz et al., 2005). These approaches suffer from noisy or incomplete trajectories. To address this, Chen et al. (2016) formulates trajectory modeling as a dictionary learning problem and uses augmented semi-nonnegative sparse coding to find motion patterns characterized by partial trajectory segments.
MoDs can also be based on velocity observations. With velocity mapping, human dynamics can be modeled through flow models. Kucner et al. (2017) presented a probabilistic framework for mapping velocity observations, named Circular-Linear Flow Field map (CLiFF-map). CLiFF-map represents local flow patterns as a multi-modal, continuous joint distribution of speed and orientation, as further described in  Sec. 3. A benefit of CLiFF-map is that it can be built from incomplete or spatially sparse velocity observations (de Almeida et al., 2024), without the need to store a long history of data or deploy advanced tracking algorithms. CLiFF-maps are typically built offline, for the reason of high computational costs associated with the building process. This constraint limits their applicability in real environments.
@@ 3.2 Learning Continuous Motion Fields
To capture temporal variations in human motion, we condition the MoD on a continuous temporal variable. Incorporating time into the input enables the model to capture temporal variations without requiring sequential rollouts, and enables efficient queries of motion dynamics at arbitrary spatio-temporal coordinates $(x,y,t)$.
@@
where $\mathbf{x}_{i}\in\mathbb{R}^{2}$ is the spatial coordinate, $\mathbf{t}_{i}\in\mathbb{R}$ is a continuous temporal variable, and $\mathbf{v}_{i}=[\rho_{i},\theta_{i}]^{\top}$ is observed velocity, we learn a continuous function $\Phi_{\theta}$ that maps a spatio-temporal coordinate $(\mathbf{x},t)$ to SWGMM parameters:
@@
$p(\mathbf{v}_{i}\mid\Phi_{\theta}(\mathbf{x}_{i},t_{i}))=\sum_{j=1}^{J}w_{j}(\mathbf{x}_{i},t_{i})\,\mathcal{N}^{\mathrm{SW}}_{\bm{\mu}_{j}(\mathbf{x}_{i},t_{i}),\,\bm{\Sigma}_{j}(\mathbf{x}_{i},t_{i})}(\mathbf{v}_{i}),$
@@
$\mathcal{L}(\theta)=-\frac{1}{N}\sum_{i=1}^{N}\log\,p\big(\mathbf{v}_{i}\mid\Phi_{\theta}(\mathbf{x}_{i},t_{i})\big).$
@@ 4 Results
To evaluate spatio-temporal maps of dynamics that capture changes of human motion patterns over time, it is essential to use datasets that span multiple days and reflect variations in human motion patterns. Our experiments were conducted using two real-world datasets, ATC (Brščić et al., 2013) and ETH (Pellegrini et al., 2009)/UCY (Lerner et al., 2007). The ATC dataset provides sufficient multi-day coverage for evaluation, consisting of long-term real-world people tracking data collected in a shopping mall. The ETH/UCY dataset contains pedestrian trajectories captured in outdoor environments such as university campuses, with multiple unique scenes. Details are in Appendix B.
@@ 4.1 Baselines
Circular-Linear Flow Field Map (CLiFF-map).
@@
Online CLiFF-map.
Online CLiFF-map (Zhu et al. (2025)) extends the static CLiFF model by updating the SWGMM parameters incrementally as new motion observations become available. Each grid location maintains an SWGMM, which is initialized upon first receiving observations and subsequently updated using the stochastic expectation-maximization (sEM) algorithm (Cappé & Moulines (2009)). In sEM, the expectation step of the original EM algorithm is replaced by a stochastic approximation step, while the maximization step remains unchanged. Like the static CLiFF-map, Online CLiFF-map outputs SWGMM parameters at each grid location, but supports continuous adaptation over time. In the experiments, we follow a spatio-temporal setting by generating an online CLiFF-map for each hour. Observations collected in an hour interval are treated as the new data batch for updating SWGMMs, producing a temporally adaptive motion representation.
Spatio-Temporal Flow Map (STeF-map).
@@ 4.2 Quantitative Results
To quantitatively evaluate the accuracy of modeling human motion patterns (MoD quality), we use the negative log-likelihood (NLL). An MoD represents human motion as a probability distribution over velocity conditioned on a spatio-temporal coordinate $(x,y,t)$, implemented as either an SWGMM (our method and CLiFF-maps) or a histogram (STeF-map). To evaluate representation accuracy, we use test data consisting of observed human motions in the same environment. For each test sample $(x,y,t)$, we query the MoD to obtain the predicted distribution and compute the likelihood of the observed motion under this distribution. A higher likelihood indicates that the predicted distribution better aligns with the observed data. We report NLL for numerical stability and easy comparison, so lower NLL values correspond to more accurate motion representations, i.e., higher quality MoDs. Table 1 and Table 2 report the accuracy results for the ATC dataset and the ETH/UCY dataset, respectively. Our method achieves the lowest NLL, outperforming all baselines. Online CLiFF-map, CLiFF-map, and STeF-map exhibit higher NLLs, with paired t-tests showing $p<0.001$ under the null hypothesis that baseline NLL is less than or equal to ours.
@@
We also compare the map building time as shown in Table 3. For the baselines, the training time corresponds to convergence on all grid cells, while for our method it corresponds to the neural network training time. Our method is computationally efficient, training on a full day of data in under 20 minutes, while achieving higher accuracy. These results highlight the practicality of continuous MoDs for real-time applications, combining both accuracy and efficiency.
@@
Table 2: Accuracy evaluation on the ETH/UCY dataset using NLL, where lower values indicate better accuracy. We report mean $\pm$ standard deviation, together with the reduction in NLL relative to our method and the corresponding 95% confidence interval (CI) of this reduction.
Method ETH HOTEL UNIV ZARA
Ours -0.384 $\pm$ 2.051 -0.838 $\pm$ 4.043 0.404 $\pm$ 1.902 -0.342 $\pm$ 2.152
CLiFF-map 0.112 $\pm$ 4.005 0.701 $\pm$ 4.533 0.518 $\pm$ 2.125 0.068 $\pm$ 4.265
Reduction vs Ours +0.496 +1.539 +0.114 +0.410
95% CI [0.344, 0.648] [1.287, 1.791] [0.049, 0.178] [0.215, 0.604]
Online CLiFF-map 0.086 $\pm$ 4.451 1.241 $\pm$ 7.142 0.577 $\pm$ 2.548 0.186 $\pm$ 5.477
Reduction vs Ours +0.470 +2.079 +0.173 +0.528
95% CI [0.307, 0.633] [1.744, 2.412] [0.098, 0.247] [0.264, 0.792]
STeF-map 2.315 $\pm$ 6.016 3.349 $\pm$ 7.660 10.932 $\pm$ 12.771 2.784 $\pm$ 7.014
Reduction vs Ours +2.699 +4.187 +10.528 +3.126
95% CI [2.472, 2.926] [3.820, 4.554] [10.187, 10.868] [2.779, 3.472]
@@ 4.3 Qualitative Results
Fig. 3 and Fig. 4 present qualitative examples of NeMo-map compared with baseline MoDs on the ATC dataset. NeMo-map captures multimodal human motion patterns, which is important in open spaces with intersecting flows. Fig. 3 presents the central open area of the ATC shopping mall. In the region highlighted by a black circle, a dominant horizontal flow intersects with a vertical flow. NeMo-map preserves both motion modes, maintaining the full continuity of the crossing traffic. In contrast, CLiFF-map fails to capture this multimodality consistently. It shows the vertical flow only at the top and bottom of the marked region, while cutting off the middle section, resulting in an incomplete representation of the crossing motion. As a result, NeMo-map achieves lower NLL values and provides accurate representation, as shown in the NLL heatmaps in the bottom row.
@@
Fig. 4 shows the NeMo-map and baseline MoDs, together with corresponding NLL heatmaps, in the east corridor of the ATC shopping mall across different times of the day. In this corridor (right side of the ATC map), human motion patterns vary throughout the day. The dominant flow is directed left/upwards in the morning, shifts direction around noon, and turns into right/downwards in the evening. The flow fields generated by NeMo-map capture such temporal variations and implicitly align with the environment’s topology, even though no explicit map was provided during training. For instance, pedestrian speeds decrease near resting benches, motion flows pass through exits, and flows follow the corridors. Same to the middle-region observations in Fig. 3, NeMo-map not only captures the dominant following flow but also preserves the flow that crosses the corridor, which baseline MoDs fail to represent consistently. The richer representation leads to lower NLL values for NeMo-map, demonstrating its higher accuracy in modeling time-varying human dynamics.
Fig. 5 presents generated MoDs in ETH/UCY datasets. The first row shows the ZARA scene at frame 600, along with the pedestrian trajectories for reference. The dominant motion trend in ZARA is along the horizontal direction. While baseline MoDs mostly conform to this dominant flow, NeMo-map adapts to an emerging upper-left directional trend. The second row shows NeMo-map results for three other ETH/UCY scenes: ETH, HOTEL, and UNIV. Pedestrian trajectories from the entire sequences are shown. NeMo-map captures diverse motion patterns. For example in the HOTEL scene, pedestrians tend to keep right when encountering oncoming flows.
Refer to captionNeMo-map
Refer to captionOnline-map
Refer to captionCLiFF-map
Refer to captionSTeF-map
Figure 3: NeMo-map and baseline MoDs in the central open area of the ATC shopping mall. NeMo-map is continuous in space, but for visualization, we query it on a uniform grid, while baseline MoDs are defined only on discrete grid cells. The top row shows the velocity distributions generated by each MoD. At each location, arrow color encodes motion orientation and arrow length encodes speed. The black circle highlights a region where horizontal and vertical flows intersect. NeMo-map preserves both motion modes, while baseline MoDs break the vertical flow continuity. The bottom row shows the corresponding NLL heatmaps, where lighter values indicate lower NLL, i.e., more accurate representations.
$09:0012:0015:0018:00$
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption NeMo-map
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption Online-map
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption CLiFF-map
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption STeF-map
Figure 4: NeMo-map and baseline MoDs in the east corridor, located on the right side of the ATC map (Fig. 7), at different times of the day. The left four columns show the generated velocity fields. Nemo-map captures the temporal variation of pedestrian flows. The right four columns show the corresponding NLL heatmaps. Compared with the baselines, NeMo-map preserves both the dominant corridor following flow and the crossing flow, leading to consistently lower NLL values.
Refer to captionZARA
Refer to captionNeMo-map
Refer to captionOnline-map
Refer to captionCLiFF-map
Refer to captionSTeF-map
Refer to captionETH
Refer to captionHOTEL
Refer to captionUNIV
Figure 5: Top row: NeMo-map and baseline MoDs in the ZARA scene at frame 600. Pedestrian trajectories are shown as reference. NeMo-map adapts to the changes from dominant horizontal flow to an emerging flow toward the upper-left entrance, while baseline MoDs remain aligned with the horizontal movement. Bottom row: NeMo-map results for the ETH, HOTEL and UNIV scenes with trajectories from the entire sequences. It captures multimodal human motion dynamics across different environments. For example, in the HOTEL scene, pedestrians tend to keep right when encountering oncoming flows.
4.4 Downstream task validation
We evaluate NeMo-map in the downstream task of long-term human motion prediction (LHMP). Accurate LHMP is essential for various applications, include optimized motion planning, advanced automated driving, and improved human-robot interaction. While short-term predictions can often rely on current state and immediate interactions, long-term predictions demand more comprehensive modeling of human motion patterns. Maps of Dynamics (MoDs) provide a way to encode such patterns and can act as a prior that guides motion prediction. Experiments of LHMP are conducted on ATC using the same testing split as in Sec. 4.2. The prediction horizon is up to $60\text{\,}\mathrm{s}$. We compare NeMo-map with MoD-based baselines: CLiFF-LHMP (Zhu et al., 2023) and STeF-LHMP (Molina et al., 2022), a diffusion-based model (Gu et al., 2022) and a transformer-based model (Shi et al., 2023). Implementation details and evaluation metrics are provided in Appendix D.
Table 4 presents LHMP performance on the ATC dataset with the prediction horizon of $60\text{\,}\mathrm{s}$. Results for other prediction horizons from $10\text{\,}\mathrm{s}$ to $60\text{\,}\mathrm{s}$ are detailed in Fig. 6. NeMo-map achieves the best performance among MoD-based predictors, with paired t-tests showing $p<0.01$, under the null hypothesis that our method yields equal or higher mean error than the baseline.
The improvements arise from NeMo-map’s ability to provide smooth and accurate motion fields across continuous space and time, in contrast to CLiFF-LHMP, which relies on a grid of SWGMMs and search for nearby cells during prediction. By modeling the velocity distribution continuously, NeMo-map captures more precise local flow patterns that result in more accurate long-term predictions. In addition, unlike STeF-map which represents orientation using an 8-bin histogram and ignores speed, NeMo-map jointly models both speed and orientation in a continuous manner, offering a more expressive motion prior. A qualitative example for a $60\text{\,}\mathrm{s}$ prediction is shown in Fig. 6. Human motion prediction using NeMo-map generates realistic trajectories that follow the complex topology of the environment, while deep learning-based results are unfeasible by crossing the walls.
Table 4: Long-term human motion prediction results on the ATC dataset with a prediction horizon of $60\text{\,}\mathrm{s}$. We report ADE/FDE (mean $\pm$ one standard deviation), where lower values indicate better performance.
Metric Ours CLiFF-LHMP STeF-LHMP TUTR MID
ADE $\downarrow$ 5.08 $\pm$ 4.28 5.45 $\pm$ 4.54 5.88 $\pm$ 5.55 12.10 $\pm$ 8.20 21.20 $\pm$ 5.28
FDE $\downarrow$ 10.97 $\pm$ 9.92 11.45 $\pm$ 10.20 12.24 $\pm$ 12.16 27.26 $\pm$ 19.23 44.47 $\pm$ 10.91
[…]